OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 2x - 9y = 1 x - 4y = 1 (1, 5) (-5, -1) (5, 1) (-5, 1)

OpenStudy (anonymous):

From the 2nd equation x=1+4y Plug this in the first one 2(1+4y)-9y=1 y=1 x=1+4y=5 so the point is (5, 1) x comes first